ELECTRONIC STRUCTURE OF GROUND AND SOME LOW-LYING EXCITED STATES OF CoCO: AN AB INITIO MOLECULAR ORBITAL STUDY

\underline{TOSHIHIDE HIKIDA}, KEIICHI TANAKA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Fukuoka 812-8581, Japan; SACHIKO S. ITONO and TSU-NEO HIRANO, Department of Chemistry, Faculty of Science, Ochanomizu University, Tokyo 112-8610, Japan; UMPEI NAGASHIMA, Tsukuba Advanced Computing Center, National Institute of Advanced Industrial Science and Technology, Ibaraki 305-8562, Japan.

MR-SDCI + Q and MR-ACPF molecular orbital calculations with the Roos-ANO basis sets have been carried out for the ground and some low-lying excited states of CoCO in order to determine the molecular constants and electronic structures, which will be comprehensive to those from the IR and millimeter-wave spectroscopies.^{a} Dynamical electron correlations for the 8-10σ electrons are taking into account in the MR-SDCI + Q calculations with the active space consisting of Co 3d, 4s orbitals and CO π, π^{∗} orbitals. Relativistic energy corrections and spin-orbit interaction were also calculated. The electronic ground state was confirmed to be ^{2}Δ_{i}, and calculated rotational constant and vibrational frequencies agreed well with experimental values. Dipole moment and spin-orbit coupling constant were predicted to be 3.8 D and -486 cm^{−1}, respectively. Energy levels and electronic structure of some low-lying excited states have also been calculated.

aHikida, et al.: two papers submitted to the session for Radicals and Ions of this 57th Symposium (2002).